# Self Supervised Networks for Learning Latent Space Representations of Human Body Scans and Motions

### v1

@@ 4.2 Generative modeling in the space of human body shapes:
Figure 10: Interpolations between 4D data: we display two interpolation examples between 4D data using our models. The top and bottom rows correspond to two 4D motion sequences with the middle rows corresponding to the interpolation based on our approach. The example on the right demonstrates the ability of the algorithm to interpolate between motions that are performed at different speeds: we show the interpolation between the first half of a sequence to the full sequence, i.e., the second sequence performs the motion in half the time. Instead of blending the motion at each fixed time our algorithm correctly extends the length of the motion in the interpolation. The highlighted shape shows the moving endpoint of the first sequence.
4.3 Interpolation and extrapolation of 4D-data
@@
As a third example, we show how our frameworks can be combined to lead to an algorithm for 4D-data interpolation (and extrapolation), i.e., we aim to interpolate (extrapolate, resp.) between two motion sequences $q_{1}(t)$ and $q_{2}(t)$. A simple algorithm for achieving this goal would consist of interpolating for each fixed time $t_{0}$ in the motion sequences between their corresponding latent codes. This can lead, however, to sub-optimal results as humans perform motions at varying speed and thus $q_{0}(t_{0})$ might not correspond to $q_{1}(t_{0})$ but rather to a different time point $q_{1}(t_{1})$. To address this challenge we resort to MoGeN, which maps motion sequences to straight lines in the lifted latent space. This allows to reduce the 4D-interpolation problem to the task of averaging lines in $\mathbb{R}^{N}$. We present two examples of this procedure in Figure 10, where the time independence of the 4D-interpolation algorithm is highlighted by the example on the right.
4.4 Representation of scans obtained from neural radiance field reconstruction
An important aspect of the proposed methods is the use of mesh invariant feature representations as inputs of the neural networks which a priori allows them to be applied to raw unregistered and noisy meshes with limited pre-processing needed beyond usual rigid alignment. This includes meshes obtained from different scanning modalities as the result of Figure 11 shows with a surface obtained from neural radiance field reconstruction that the authors directly acquired using their cellphone camera. One shortcoming this result points to, however, is the oversimplification of the body shape for subjects too different from those that can be represented in the latent space. We should note that this is primarily a limitation of the latent space used for body shape representation. As the generic framework of this paper is not tied to a specific choice of latent representation, we hope to address such issues in future iterations of this work.
@@ 4.4 Representation of scans obtained from neural radiance field reconstruction
Acknoledgements
N.C and E.H were supported by the National Science Foundation, under the grants DMS-2438562 and DMS-2402555. M. B. was supported by the National Science Foundation, under the grant DMS-2324962, and by the Binational Science Foundation (BSF).
Appendix A Ablation Study using a different Latent space models
A.1 The VariShape model using the elastic shape analysis latent space of [15]
In Section 4, we reported the performance of VariShaPE using SMPL [19] as the latent model $F$ and the model proposed by [15] as the feature map $G$. In Table 4, we further explore the performance of the VariShaPE for alternate combinations of $F$ and $G$. These combinations allow us to examine the robustness and adaptability of the VariShaPE framework when employing different underlying latent space and feature models.
F: BaRe[15]   G: BaRe[15]
Registered Unregistered
Mean Error 2.1mm NA
Varifold Error 0.034 0.037
Chamfer Error 2.5mm 3.0mm
F: SMPL[19]   G: SMPL[19]
Registered Unregistered
Mean Error 2.2mm NA
Varifold Error 0.037 0.041
Chamfer Error 2.8mm 3.4mm
Table 4: Ablation study for various choices of $F$ and $G$ in VariShaPE. In the left table, we present results with $F$ and $G$ chosen to be the latent space model of [15]. The right table contains the results where $F$ and $G$ are chosen to be the latent space model of SMPL [19].
The VariShaPE model achieves consistently strong performance across these variations, demonstrating its effectiveness regardless of the specific models used for $F$ and $G$. Nonetheless, the combination highlighted in the main paper yields slightly superior results, suggesting it is the optimal configuration among those tested.
A.2 The MoGen model for the elastic shape analysis latent space of [15]
Section 4 also reports and discusses the results of the MoGeN model applied to the SMPL latent space [19]. To further investigate the model’s versatility, Table 5 presents the results of MoGeN applied to the latent space defined by [15]. These results indicate that it performs comparably well within this alternative latent space.
@@ A.2 The MoGen model for the elastic shape analysis latent space of [15]
Table 5: Interpolation and Extrapolation Results for the model of [15]: We present the mean vertex error and mean varifold error of paths in the latent space of [15] produced by the MoGeN model. We report the results for the testing set of mini-sequences extracted from the DFAUST testing set.
The findings suggest that the MoGeN model maintains its effectiveness regardless of the choice of latent space, performing similarly to its results with the SMPL model.
Appendix B Ablation Study on the Choice of Loss Function in VariShaPE
Finally, we conduct an ablation study on the choice of loss function used in training of the VariShaPE model. We present the results for the model trained with these losses evaluated on an unseen testing set.
Registered Unregistered
Mean Vertex Dist. Varifold Chamfer Varifold Chamfer
Mean Vertex 2.1mm 0.035 2.5mm 0.036 2.9mm
Chamfer 2.0mm 0.030 2.0mm 0.035 3.3mm
Varifold 2.4mm 0.028 2.7mm 0.033 3.6mm
Varifold + Mean Vertex 1.8mm 0.026 2.2mm 0.033 3.4mm
Varifold + Chamfer 2.1mm 0.030 2.2mm 0.036 3.0mm
Table 6: Ablation Study for training the VariShaPE model with different choices of loss function.
While the combination of mean vertex error and varifold error yields marginally superior results, we focus on the model trained exclusively with mean vertex error. This decision is motivated by practical considerations, as the mean vertex error-only model achieves nearly equivalent performance while reducing training time by approximately 4 hours.

### v2

@@ 4.1 Mapping Registered and Unregistered Scans to SMPL
Figure 8: Interpolations between 4D data: We display two examples interpolations between 4D data using our models. The top and bottom rows correspond to motions from 4D data with the middle rows corresponding to the interpolation using our method. The example on the right demonstrates the capability of the algorithm to interpolate between motions, that are performed with a different speed: we show the interpolation between the first half of a sequence to the full sequence, i.e., the second sequence performs the motion in half the time. Instead of blending the motion at each fixed time our algorithm correctly extends the length of the motion in the interpolation. The highlighted shape shows the moving endpoint of the first sequence.
4.2 Motion Interpolation and Extrapolation:
Next we investigate the capabilities of our second proposed network, MoGeN, to accurately learn the geometry of a latent space representations of human body motions. Therefore, we compare our method’s ability to (1) predict human motions interpolating between a start and end point and (2) to extrapolate upon initial motions. To evaluate the performance of MoGeN we compare our results again to two competing methods: first, we compare to linear interpolations in the pose parameter space of the SMPL model. This is a natural baseline to compare our model against and allows us to visualize comparisons of our model in the latent space, cf. Figure 5 and Figure 11 in Appendix B. Secondly, we provide quantitative comparisons of the interpolations and extrapolations produced by our method and the auto-decoder latent space construction of [9]. In Tab. 2 one can observe that the proposed algorithm outperforms the competing methods, which is also visualized in the qualitative example in Figure  5 and Figure 11.
4.3 Motion Transfer, Generative Modeling and Interpolation between 4D-data:
Having demonstrated the accuracy of the proposed network architectures, we aim to present three different applications of the proposed methodology. The first one consists of a real-time approach for motion transfer. Fig. 6 shows two different examples and an explanation of the procedure that uses VariShaPE to encode the given sequence and target in latent space, and subsequently employs MoGeN to accurately transport the corresponding path of latent codes to the latent codes of the target shape.
Next, we demonstrate how one can combine MoGeN and VariShaPE to obtain a generative model for human bodies based on estimating an empirical distribution from observed training data, cf. Fig. 7 for more details on the procedure. In that same figure, we show 18 random samples, where the empirical distribution was constructed using 30,000 scans. Note that this procedure is only feasible due to the light speed encoding capabilities for raw scans provided by VariShaPE: for our experiment the encoding of the whole population of unregistered training scans took less than one minute.
@@ 4.3 Motion Transfer, Generative Modeling and Interpolation between 4D-data:
Finally, we show how our frameworks can be combined to lead to an algorithm for 4D-data interpolation (and extrapolation), i.e., we aim to interpolate (extrapolate, resp.) between two motion sequences $q_{1}(t)$ and $q_{2}(t)$. A simple algorithm for achieving this goal would consist of interpolating for each fixed time $t_{0}$ in the motion sequences between their corresponding latent codes. This can lead, however, to suboptimal results as humans perform motions at varying speed and thus $q_{0}(t_{0})$ might not correspond to $q_{1}(t_{0})$ but rather to a different time point $q_{1}(t_{1})$. To address this challenge we resort to MoGeN, which maps motion sequences to straight lines in the lifted latent space, which reduces the 4D-interpolation problem to the task of averaging lines in $\mathbb{R}^{N}$. We present two examples of this procedure in Figure 8, where the time independence of the 4D-interpolation algorithm is highlighted in the example on the right.
Registered Unregisterd
Cham. Ours Cham. Ours
Mean Error 0.083 0.021 NA NA
Varifold Error 0.054 0.034 0.047 0.037
Chamfer Error 0.026 0.025 0.032 0.030
Registered Unregisterd
Cham. Ours Cham. Ours
Mean Error 0.063 0.022 NA NA
Varifold Error 0.047 0.037 0.048 0.041
Chamfer Error 0.031 0.028 0.057 0.034
Table 3: Ablation Study for Choices of Latent Space Models in the VariShaPE Model: In the left table, we present results for the VariShaPE model, with $F$ and $G$ chosen to be the latent space model of [5, 6]. The right table contains the results for the VariShaPE model, with $F$ and $G$ chosen to be the latent space model of SMPL [1].
@@
4.4 Ablation Study using a different Latent Space Model
In the previous part we reported the results of the VariShape model with $F$ as the SMPL model [1] and $G$ as the model of [5, 6]. In Table 3 we present the experimental results for other combinations of choices for $F$ and $G$. Notably, the performance of the VariShaPE model for these choices is similar to the one presented previously. Moreover, we reported and discussed the results of the MoGeN model using the SMPL latent space. In Table 4, we present the results of the MoGeN model where the latent space is instead that of [5, 6].
@@ 4.4 Ablation Study using a different Latent Space Model
Table 4: Interpolation and Extrapolation Results for the model of [5, 6]: We present the mean vertex error and mean varifold error of paths in this latent space as produced by the MoGeN model. The reported results are for a test set of mini-sequences extracted from the DFAUST testing set.
5 Conclusions, Limitations and Future Work
In summary, we have proposed in this paper new self-supervised models for learning not only latent space representations of human body shape and pose but also the underlying geometry of the latent space itself. The latter is achieved through the introduction of an additional lifted space in which observed body pose motions get unfolded to straight paths. We have further shown how these two frameworks can be used to perform common tasks such as motion interpolation, extrapolation and transfer, and to estimate generative models in the space of human body pose and shape.
An important aspect of the proposed methods is the use of mesh invariant feature representations as inputs of the neural networks which a priori allows them to be applied to raw unregistered and noisy meshes with limited pre-processing needed beyond usual rigid alignment. This includes meshes obtained from different scanning modalities as the result of Fig. 9 shows with a surface obtained from neural radiance field reconstruction that the authors directly acquired using the camera of their phone. One shortcoming this results points to, however, is the oversimplification of the body shape for subjects too different from those that can be represented in the latent space. We should note that this is primarily a limitation of the latent space used for body shape representation. As the generic framework of this paper is not tied to a specific choice of latent representation, we hope to address such issues in future iterations of this work.
Appendix A Qualitative Comparison to [9] for Registered Data.
Table 1 of the main article reported the quantitative comparison of our model for both registered and unregistered meshes. We further complement those results by emphasizing, in Figure 10, the qualitative differences in reconstruction on registered meshes of our proposed method and the variational autoencoder framework described in [9].
Figure 10: Qualitative Comparison of SMPL latent code retrieval methods for registered data: We display 7 registered human meshes (middle), reconstructions produced by the VAE (top), and reconstructions produced via our model (bottom).
Appendix B Qualitative Comparison of MoGeN for Extrapolations.
In Figure 5 of the main paper, we display the qualitative comparison of MoGeN with linear interpolations in SMPL space. In Figure 11 we display a similar comparison of MoGeN with linear extrapolations in SMPL space. This example highlights the strength of our model for learning complex human motions.
Figure 11: Qualitative comparison of MoGeN with linear extrapolation in SMPL space. In red, we display a sequence of meshes corresponding to a human motion from data with the initial points highlighted. In green, we display the path which arises from linear extrapolation from the initial points in the latent space. Finally, in blue, we display the path produced by the MoGeN extrapolation model with the boundary points as inputs. On the bottom we visualize these paths in two coordinates of the latent space with the same color coding.